# Firstname1 Lastname1    Firstname2 Lastname2    Firstname3 Lastname3    Firstname4 Lastname4    Firstname5 Lastname5    Firstname6 Lastname6    Firstname7 Lastname7    Firstname8 Lastname8    Firstname8 Lastname8

### v1

@@ Abstract
Transformers excel at discovering patterns in sequential data, but understanding their limits and how they learn remains a crucial topic. In this paper, we investigate the ability of Transformers to learn pseudo-random number sequences from linear congruential generators (LCGs), defined by the recurrence relation $x_{n+1}=ax_{n}+c\;\mathrm{mod}\;m$. Our analysis reveals that with sufficient architectural depth, context length, and training data variety, Transformers can predict LCG sequences with unseen moduli ($m$) and parameters ($a,c$). We demonstrate successful learning up to $m=2^{32}$, and we perform systematic analyses of the effect of dataset complexity, architecture size, and context lengths. We uncover emergent structures in the embedding layers and attention heads, which we use to reverse engineer the underlying algorithms used by the Transformer. We first consider training and testing on a single modulus, which reveals that models learn to factorize the modulus and subsequently utilize digit-wise number representations to make sequential predictions. When generalizing to unseen moduli, the models employ a greedy approach to estimate the unknown modulus from the context and then use prime factorizations to generate predictions. In the latter case, we observe a transition at a critical depth $d=3$, below which the Transformer is incapable of learning. We also find that the number of in-context sequence elements needed to reach high accuracy scales sublinearly with the period.
@@ 1 Introduction
Transformer-based language models have proven to be extremely powerful sequence generative models. With copious amounts of training data and computational resources, they can identify and learn complex patterns from training corpora, resulting in numerous remarkable capabilities (attn_is_all_you_need; dosovitskiy2021an). Recent research has demonstrated that these models, when provided with sufficient context and inference compute, can acquire new patterns and capabilities without additional training through techniques such as in-context learning (radford2019language) and chain-of-thought reasoning (wei2023chain). While these models have achieved unprecedented success, understanding what underlying patterns are learned and how they learn them remains a significant challenge.
Pseudo-Random Number Generators (PRNGs) represent an interesting test case for exploring these challenges. These algorithms, which are fundamental to modern cryptography and computer science, are designed to produce outputs that pass statistical tests for randomness, but nevertheless arise from mathematical patterns that could potentially be learned by sufficiently powerful sequence models.
This intersection between Transformer models’ pattern-learning capabilities and the structured nature of PRNG outputs raises intriguing questions about both the capabilities and limitations of these models. Can Transformers learn to predict PRNG outputs given sufficient training data, model capacity and context? If so, what implications does this have for our understanding of both Transformer architectures and PRNGs? Do the Transformers learn the underlying generating algorithm or merely detect shortcuts and spurious patterns? What effect do model capacity, data variety, training methodologies and context length have on the capabilities of Transformers?
This work aims to answer these questions by focusing on learning sequences obtained from linear congruential generators (LCGs) using GPT-style autoregressive Transformers. We demonstrate how Transformers can successfully learn LCGs and training with moduli up to $m=2^{32}$. We perform interpretability analyses, uncovering emergent structures in the embedding layers, attention heads, and underlying algorithms that the Transformer uses to learn the sequences. We also perform several systematic scaling analyses to understand the effect of architecture and sequence complexity on model performance and in-context learning ability.
@@ 1.1 Related works
Interpretability and Modular Arithmetic A growing body of work uncovers the circuits, algorithms, and emergent structures learned by Transformers (sharkey2025interp; olsson2022context; Ahn2023gradient; vonoswald2023Transformers; akyurek2023what; hendel2023incontext; liu2024incontextvector). A particularly fruitful setting involves simple problems in modular arithmetic (power2022grokking; gromov2022grokking; nanda2023progress; zhong2023clock; doshi2024to; he2024learning). Our work adds to this by reverse-engineering the underlying algorithms and uncovering emergent structures in learning pseudo-random number sequences.
Cracking PRNGs: There is a classic duality between cryptography and learning theory (rivest1991cryptography), and cracking PRNGs is an important topic in cryptography. Nevertheless, deep learning-based attacks have received limited attention in the post-Transformer era. amigo2021 demonstrated that a fully-connected neural network can predict the outputs of a modified LCG with fixed (irrational) parameters $(a,c,m)=(1,\pi,1)$. In comparison, we systematically analyze the harder cases of unseen parameters using Transformers, reverse-engineer the learned algorithms, and study effects of scale and complexity.
Context-Free Grammar: LCG can also be viewed as a formal language (Type-3 regular grammar) lying within the Chomsky hierarchy (chomsky1956three). Formal languages provide an interesting setting for synthetic datasets that can be used to understand the properties of neural networks in controlled settings (deletang2023chomsky; allenzhu2024physicslanguagemodels1; cagnetta2024deep; cagnetta2024towards).
Chaotic time-series: A major application of neural networks is predicting time-series for chaotic dynamics, such as weather prediction (lam2023learning) and financial modeling. PRNGs provide an analog of such dynamics in the discrete setting.
@@ 1.2 Linear Congruential Generators
LCG is a simple PRNG that generates the next number in a sequence $\{x_{n}\}_{n=0}^{N}$ according to the linear map:
$x_{n+1}=(ax_{n}+c)\mod m,$ (1)
where $m>0$ is the modulus, $0<a<m$ is the multiplier and $0\leq c<m$ is referred to as the increment. An LCG map is uniquely defined by the choice of $m,a,c$ and the initial seed $x_{0}$. An important quantity that determines the complexity of an LCG sequence is its period: $1\leq\mathcal{T}_{m}(a,c)\leq m$. As we will show in the following sections, the period of a sequence plays a major role in the difficulty of prediction with Transformers. According to the Hull-Dobell Theorem (hull-dobell), the period $\mathcal{T}_{m}(a,c)=m$ if and only if the values of $a$ and $c$ satisfy the following criteria:
  1. I.
    $m$ and $c$ are coprime
  2. II.
    $a-1$ is divisible by all prime factors of $m$
  3. III
    $a-1$ is divisible by $4$ if $m$ is divisible by $4$.
We will evaluate all our models exclusively on sequences that obey the criteria of this theorem.
LCGs are widely utilized for their speed and simplicity, often forming the core of more complex PRNGs like PCG-64, which is used in NumPy. LCGs perform poorly at small bit sizes but improve rapidly with larger state sizes. For instance, an LCG with 88 bits of state can pass the stringent BigCrush randomness test (ONeill2014PCGA).
@@ 2.1 Dataset Generation and Evaluation
We investigate two training paradigms: Fixed Modulus (FM) and Generalization to Unseen Modulus (UM).
Fixed Modulus (FM): We first apply the Hull–Dobell Theorem to determine possible values of $(a,c)$ that maximize the sequence period for a given modulus $m$. From these possible values, we randomly sample $64$ values of each $a$ and $c$ to generate the test dataset. To generate the training dataset, we exclude these choices of $a,c$ and uniformly sample 100k LCG sequences of length $L$ (context length), with different $a,c$ and $x_{0}$ (regardless of their period).
Generalization to Unseen Modulus (UM): In this more challenging paradigm, we first select test modulus $m_{\text{test}}$ that are excluded from training. The training dataset incorporates $n_{m}$ different modulus uniformly sampled from the range $[L,m_{\text{max}}]$, where $m_{\text{max}}=\lfloor 1.2\cdot m_{\text{test}}\rfloor$. For each modulus $m$, we uniformly select $n_{a}$ multipliers ($0<a<m$) and $n_{c}$ increments ($0\leq c<m$), generating sequences of length $L$ with random initial values $x_{0}$. This results in a total of $N=n_{m}\times n_{a}\times n_{c}$ training sequences. To construct the test dataset for each $m_{\text{test}}$, we randomly sample 64 values each of $a$ and $c$ that generate maximum-period sequences.
In both paradigms, test accuracy is measured by averaging over all sequences generated from the excluded values of $a$ and $c$ and 128 random initial values $x_{0}$ for each $(a,c)$ pair.
The above settings are used in Sections 3 and 4. In order to achieve better performance, we used a larger and slightly higher-quality dataset in Section 5, which we detail later.
2.2 Tokenization, Architecture and Optimizer
We employ two different settings throughout this paper. (i) In Sections 3 and 4, we tokenize each number as a unique token, with dictionary size $m$ for the FM case and $m_{\mathrm{test}}$ for the UM case. We then use GPT-style Transformers with learnable positional embeddings and weight tying (press2017tying). (ii) In Section 5, we tokenize each number using byte-tokenization with various base choices and apply abacus positional embeddings (mcleish2024abacus)
Model architecture is characterized by number of blocks (depth), embedding dimension ($d_{\mathrm{model}}$), number of attention heads ($n_{\mathrm{heads}}$). Models are trained with AdamW optimizer.
@@ 3 Training Results
Figure 1: Training phase diagram and dynamics. (a) Test accuracy for last token of models trained with $m_{\mathrm{test}}=4096$, evaluated across various depths and $n_{\mathrm{head}}$ values, showing a minimum depth requirement of $3$ for generalization; (b) Training curves for last token of a model with depth $6$ and $n_{\mathrm{head}}=4$ ($m_{\mathrm{test}}=2048$), showing that test accuracy for last token “groks” simultaneously with performance on sequences with period longer than context length $L=256$, indicating delayed discovery of underlying rules.
In this section, we focus on the more sophisticated behavior exhibited by models trained with UM data, leaving the discussion of FM for Appendix A. We begin by investigating the minimal model that can solve the task. In Figure 1 (a), we present the test accuracy for the last token of models trained on a dataset with $m_{\mathrm{test}}=4096$, varying both model depth and number of attention heads ($n_{\mathrm{heads}}$) while keeping head dimensions constant. Our results indicate that a minimum depth of $3$ is necessary for good generalization performance, with a sharp transition upon lowering the depth. Further analysis across multiple $m_{\mathrm{test}}$ values (see Appendix B) confirms that this minimal depth requirement is universal.
@@
We then carefully examine the training dynamics of a model with $n_{\mathrm{head}}=4$ and depth $6$ in Figure 1 (b). We categorize the training sequences into two groups: i) sequences with periods shorter than the context length, which can be solved through simple copying, and ii) sequences with periods longer than the context length, which require the model to deduce underlying rules for prediction. Our analysis reveals that the model first acquires copying ability for group i) in the early stages of training, and later “groks” the solution for group ii). Notably, the model’s ability to generalize to test $m_{\mathrm{test}}$ emerges simultaneously with this grokking phenomenon. These results demonstrate that the model develops different capabilities at distinct stages of training, with generalization ability emerging only after the model learns the underlying rules through solving the more challenging sequences. In Appendix C, we present an ablation study where models are trained exclusively on either short-period or long-period sequences. Our findings indicate that training exclusively on long-period sequences enables model generalization and eliminates the grokking phenomenon.
@@ 4 Interpreting how Transformers predict PRNGs
Interpreting the underlying algorithms implemented by the models is essential for understanding the models’ training dynamics and scaling capabilities. We uncover these algorithms for both FM and UM cases. While certain details of the algorithms differ in the two cases, they share some common properties originating from the underlying LCG structures.
We first discuss properties of LCG sequences that will be useful to interpret model behaviors.
4.1 Mixed Radix Representations
@@
where $\{\alpha_{2^{1}},\dots,\alpha_{2^{11}}\}$ are the binary digits (bits), with $\alpha_{2^{j}}\in\{0,1\}$.
In this binary representation, each digit (bit) has its own period along the LCG sequence. As shown in Figure 2, for a sequence of period $\mathcal{T}_{m}=m$, the $w^{th}$ lowest digit has a period of $2^{w}$ along the sequence. Thus, lower (higher) digits have smaller (larger) period along LCG sequences.
While the binary representation is natural for moduli of the form $m=2^{w}$, it is possible to generalize this representation to the case of composite moduli. Consider sequences with composite modulus $m$, which has a prime factorization $m=p_{1}^{w_{1}}p_{2}^{w_{2}}\cdots p_{i}^{w_{i}}$. In this case, the numbers in the sequence can be represented uniquely in a mixed radix form:
$n=\sum_{w=1}^{w_{1}}\alpha_{p_{1}^{w}}\,p_{1}^{w}+\sum_{w=1}^{w_{2}}\alpha_{p_{2}^{w}}\,p_{2}^{w}+\cdots+\sum_{w=1}^{w_{i}}\alpha_{p_{i}^{w}}\,p_{i}^{w}\;;$ (3)
where the digits $\alpha_{p_{j}^{w}}\in\{0,1,\dots,p_{j}-1\}$ are the generalization of bits. When $\mathcal{T}_{m}=m$, the period of each digit $\alpha_{p_{j}^{w}}$ has a period of $p_{j}^{w}$. We will see later that trained Transformers have emergent structures that find such digit-wise representations and utilize them to make systematic predictions from the context. The per-digit period will play an important role in the prediction accuracy of that digit.
Consider the $r$-step iteration of Equation 1:
$x_{n+r}=a^{r}x_{n}+\sum_{i=1}^{r}a^{i-1}c\mod m\;.$ (4)
In this new sequence, the period of each digit $\alpha_{p_{j}^{w}}$ reduces from $p_{j}^{w}$ to $p_{j}^{w}/\gcd(r,p_{j}^{w})$. Consequently, some of the higher digits become relatively simpler to predict due to reduced periods; and some lower digits become trivial to predict due to being constant. We show an example of this for $m=2048$ in Figure 2 (top panel). The digit-wise periods in the new sequence with $r=2$ reduce by a factor of 2, while the last digit is simply constant. Higher values of $r=2^{k}$ will lead to even further simplifications of the sequence. We will see in the later subsections that the Transformer has emergent structure that leverages this simplification to accurately predict LCG sequences.
For general $m$, LCG sequences simplify when $r=p_{1}^{k_{1}}p_{2}^{k_{2}}\cdots$. This simplification from period-reduction is only visible when we represent the numbers in the sequence in their mixed radix representations corresponding to $m$. A different representation would only partially (or not at all) reduce the per-digit periods.

### v2

# Tao Tao    Darshil Doshi    Dayal Singh Kalra    Tianyu He    Maissam Barkeshli

@@ Abstract
Transformers excel at discovering patterns in sequential data, yet their fundamental limitations and learning mechanisms remain crucial topics of investigation. In this paper, we study the ability of Transformers to learn pseudo-random number sequences from linear congruential generators (LCGs), defined by the recurrence relation $x_{t+1}=ax_{t}+c\;\mathrm{mod}\;m$. We find that with sufficient architectural capacity and training data variety, Transformers can perform in-context prediction of LCG sequences with unseen moduli ($m$) and parameters ($a,c$). By analyzing the embedding layers and attention patterns, we uncover how Transformers develop algorithmic structures to learn these sequences in two scenarios of increasing complexity. First, we investigate how Transformers learn LCG sequences with unseen ($a,c$) but fixed modulus; and demonstrate successful learning up to $m=2^{32}$. We find that models learn to factorize $m$ and utilize digit-wise number representations to make sequential predictions. In the second, more challenging scenario of unseen moduli, we show that Transformers can generalize to unseen moduli up to $m_{\text{test}}=2^{16}$. In this case, the model employs a two-step strategy: first estimating the unknown modulus from the context, then utilizing prime factorizations to generate predictions. For this task, we observe a sharp transition in the accuracy at a critical depth $d=3$. We also find that the number of in-context sequence elements needed to reach high accuracy scales sublinearly with the modulus. [^2] We open source the code to reproduce our results: https://github.com/dayal-kalra/transformer-prng.git
@@ 1 Introduction
Transformer-based language models have proven to be extremely powerful sequence generative models. With copious amounts of training data and computational resources, they can identify and learn complex patterns from training corpora, resulting in numerous remarkable capabilities (Vaswani et al., 2017; Dosovitskiy et al., 2021). Recent research has demonstrated that these models, when provided with sufficient context and inference compute, can acquire new patterns and capabilities without additional training through techniques such as in-context learning (Radford et al., 2019) and chain-of-thought reasoning (Wei et al., 2023). While these models have achieved unprecedented success, understanding what underlying patterns are learned and how they learn them remains a significant challenge. Pseudo-Random Number Generators (PRNGs) represent an interesting test case for exploring these challenges. These algorithms, which are fundamental to modern cryptography and computer science, are designed to produce outputs that pass statistical tests for randomness, but nevertheless arise from mathematical patterns that could potentially be learned by sufficiently powerful sequence models.
This intersection between Transformer models’ pattern-learning capabilities and the structured yet obfuscated nature of PRNG outputs raises intriguing questions about both the capabilities and limitations of these models. Can Transformers learn to predict PRNG outputs given sufficient training data, model capacity, and context? If so, what implications does this have for our understanding of both Transformer architectures and PRNGs? Do the Transformers learn the underlying generating algorithm or merely detect shortcuts and spurious patterns? What effect do model capacity, data variety, training methodologies, and context length have on the capabilities of Transformers?
This work aims to answer these questions by focusing on learning sequences obtained from linear congruential generators (LCGs) using GPT-style autoregressive Transformers. We demonstrate how Transformers can successfully learn LCGs with moduli up to $m=2^{32}$. We perform interpretability analyses, uncovering emergent structures in the embedding layers, attention heads, and underlying algorithms that the Transformer uses to learn the sequences. We also perform several systematic scaling analyses to understand the effect of architecture and sequence complexity on model performance and in-context learning ability.
@@ 1.1 Related works
Interpretability and Modular Arithmetic: A growing body of work examines the circuits, algorithms and structures learned by Transformers (Sharkey et al., 2025; Olsson et al., 2022; Ahn et al., 2023; von Oswald et al., 2023; Akyürek et al., 2023; Hendel et al., 2023; Liu et al., 2024). A notably fruitful setting involves simple modular arithmetic problems (Power et al., 2022; Gromov, 2023; Nanda et al., 2023; Zhong et al., 2023; Doshi et al., 2024a, b; He et al., 2024). Our work adds to this by reverse-engineering the underlying algorithms and uncovering emergent structures in learning pseudo-random number sequences.
Cracking PRNGs: There is a classic duality between cryptography and learning theory (Rivest, 1991), and cracking PRNGs is an important topic in cryptography. Nevertheless, deep learning-based attacks have received limited attention in the post-Transformer era. Amigo et al. (2021) demonstrated that a fully-connected neural network can predict the outputs of a modified LCG with fixed (irrational) parameters $(a,c,m)=(1,\pi,1)$. In comparison, we systematically analyze the harder cases of unseen parameters using Transformers, reverse-engineer the learned algorithms, and study effects of scale and complexity.
Formal Grammars: LCG can also be viewed as a formal language (Type-3 regular grammar) lying within the Chomsky hierarchy (Chomsky, 1956). Formal languages provide an interesting setting for synthetic datasets that can be used to understand the properties of neural networks in controlled settings (Delétang et al., 2023; Allen-Zhu & Li, 2024; Cagnetta et al., 2024; Cagnetta & Wyart, 2024).
Chaotic time-series: A major application of neural networks is predicting time-series for chaotic dynamics, such as weather prediction (Lam et al., 2023) and financial modeling. PRNGs provide an analog of such dynamics in the discrete setting.
@@ 1.2 Linear Congruential Generators
LCG is a simple PRNG that generates the next number in a sequence $(x_{0},x_{1},\dots,x_{t})$ according to the map:
$x_{t+1}=(ax_{t}+c)\mod m,$ (1)
where $m>0$ is the modulus, $0<a<m$ is the multiplier and $0\leq c<m$ is referred to as the increment. An LCG map is uniquely defined by the choice of $m,a,c$ and the initial seed $x_{0}$. An important quantity that determines the complexity of an LCG sequence is its period: $1\leq\mathcal{T}_{m}(a,c)\leq m$. As we will show in the following sections, the period of a sequence plays a major role in the difficulty of prediction with Transformers. According to the Hull-Dobell Theorem (Hull & Dobell, 1962), the period $\mathcal{T}_{m}(a,c)=m$ if and only if the values of $a$ and $c$ satisfy the following criteria: (i) $m$ and $c$ are coprime, (ii) $a-1$ is divisible by all prime factors of $m$, (iii) $a-1$ is divisible by $4$ if $m$ is divisible by $4$. We evaluate (test) all our models exclusively on sequences that obey the criteria of this theorem.
LCGs are widely utilized for their speed and simplicity, often forming the core of more complex PRNGs like PCG-64, which is used in NumPy. LCGs perform poorly at small bit sizes but improve rapidly with larger state sizes. For instance, an LCG with 88 bits of state can pass the stringent BigCrush randomness test (O’Neill, 2014).
@@ 2 Training Setup
We train decoder-only Transformers to autoregressively predict the next number in LCG sequences. This means it takes as input an LCG sequence $(x_{0},\cdots,x_{L-1})$, outputs a sequence $(y_{0},\cdots,y_{L-1})$, and trained so $y_{t}$ matches $x_{t+1}$.
To predict an unknown LCG sequence, the Transformer needs to infer $m,a$, and $c$ in-context. We test the model’s generalization ability in two distinct paradigms of increasing difficulty: FM: The model is trained and tested on sequences with a fixed modulus $m$. UM: Model is trained on varying moduli, and tested on unseen moduli $M_{\mathrm{test}}=\{m_{\mathrm{test}}\}$. We highlight the key details of our experimental setups here and provide an extensive discussion in Appendix A.
@@ 2.1 Dataset Generation and Evaluation
The settings below are used in Sections 3 and 4. In order to achieve better performance, we used a larger and higher-quality dataset in Section 5, which we detail later.
Fixed Modulus (FM): Given a modulus $m$, we apply the Hull-Dobell Theorem to determine the possible values of ($a,c$) that maximize the period. We then randomly select $64$ values of $a$ and $c$ to generate the test dataset. To generate the training dataset, we exlude these test choices of ($a,c$) and uniformly sample $N=100,000$ LCG sequences of length $L+1$ (where $L$ is the context length), with $n_{a}$ values of multipliers and $n_{c}$ values of increments. For each set of parameters ($a,c$), we sample an LCG sequence with a randomly selected initial seed $x_{0}$. Note that the training dataset includes sequences with varying periods, while the test data only contains sequences that maximize the period.
Generalization to Unseen Modulus (UM): In this more challenging paradigm, we first select a set of test moduli $M_{\text{test}}=\{m_{\text{test}}\}$ that would be reserved exclusively for evaluation. For each test modulus $m_{\text{test}}\in M_{\text{test}}$, we determine the values of ($a,c$) that maximize the period. We then randomly select $64$ values of $a$ and $c$ each to generate the test dataset. These $64^{2}$ ($a,c$) pairs are not considered while generating the training dataset.
For the training dataset generation, we sample $n_{m}$ modulus values from the range $[L,m_{\text{max}}]$, with $m_{\text{max}}=\lfloor 1.2\max(M_{\text{test}})\rfloor$, while excluding all the values in $M_{\text{test}}$. For each modulus value $m$, we uniformly select $n_{a}$ multipliers and $n_{c}$ increments, excluding the ones reserved for testing. For each triplet ($a,c,m$), we generate a sequence of length $L+1$ using a randomly selected initial seed $x_{0}$. This results in a total of $N=n_{m}\times n_{a}\times n_{c}$ training sequences.
We found that $n_{m}\gtrsim m_{\text{test}}/4$ yields good generalization performance. Based on this relationship and our target total number of training examples $N$, we sample $n_{a}=n_{c}=\sqrt{\frac{N}{m_{\text{test}}/4}}$ values of multipliers and increments. Unless explicitly specified, we use this setting as the default configuration for all experiments.
In both paradigms, test accuracy is averaged over all $(a,c)$ test pairs and multiple initial seeds $x_{0}$ per pair. Accuracies are tracked at all sequence positions $1\leq t\leq L$.
2.2 Tokenization, Architecture, and Optimizer
In Sections 3 and 4, each number is tokenized as a unique token, using a dictionary of size $m$ (FM) or $m_{\mathrm{max}}$ (UM). We employ GPT-style Transformers with learnable positional embeddings and weight tying (Press & Wolf, 2017).
When we scale up to larger moduli in Section 5, we restrict the dictionary size to $b$ by tokenizing each number in the sequence in base-$b$ (e.g. $b=2^{8},3^{5}$). This results in $\lceil\log_{b}{m}\rceil$ tokens for each number. We also apply (modified) abacus positional embeddings (McLeish et al., 2024).
The model architecture is characterized by the number of blocks (depth), embedding dimension ($d_{\mathrm{model}}$), number of attention heads ($n_{\mathrm{heads}}$). Models are trained with AdamW (Loshchilov & Hutter, 2019) and CrossEntropy loss.
@@ 3 Training Results
Figure 1: Accuracy of predicting the last number (token) in the sequence: phase diagrams w.r.t. various depths and $n_{\mathrm{heads}}$ values. (a) $m_{\mathrm{test}}=2048$, constant width: $d_{\mathrm{model}}=768$. (b) $m_{\mathrm{test}}=4096$, width scaled proportionally: $d_{\mathrm{model}}=128\times n_{\mathrm{heads}}$.
We begin by investigating the minimal model that can solve the two tasks in consideration. Surprisingly, we found that Transformers only require one layer and one attention head to learn the FM task, as shown in Figure 2 (a) (for further results, see Appendix B). Conversely, the UM task requires a stronger architecture and careful hyperparameter tuning. Appendix C shows that model performance depends on the modulus, with prime moduli being challenging in the FM setting but not in the UM setting.
In Figure 1 we show how the performance varies with model depth and the number of attention heads. In Figure 1(a) we keep the embedding dimension fixed to $d_{\text{model}}=768$, whereas in Figure 1(b) we scale it proportionally to the number of heads ($d_{\text{model}}=128\times n_{\textrm{heads}}$.). In both cases, we find that a minimum of three layers are required for effective generalization, with performance degrading sharply below this threshold. Further analysis across multiple $m_{\mathrm{test}}$ values (see Appendix D) confirms that this minimal depth requirement is universal. We also observe that additional attention heads improve model performance, with substantial gain occurring when increasing from one to two heads.
Several prior studies have also observed sharp changes in model capabilities as a function of model depth. This includes induction head formation (Olsson et al., 2022), in-context learning of modular addition (He et al., 2024) and various in-context generalization tasks (Chen & Zou, 2024). In general it is unclear to what extent these sharp depth-dependences are due to jumps in expressivity or trainability.
The UM task shows strong sensitivity to hyperparameters. As we increase $m_{\text{test}}$ while keeping the model size fixed, we observe two key phenomena: the optimal learning rate ($\eta$) and weight decay strength ($\lambda$) shift significantly, and simultaneously, the range of hyperparameter resulting in effective performance narrows (see Section A.5).
@@
Figure 2: Training/test accuracy curves for predicting last number (token). (a) FM: ($m=2048,\mathrm{depth}=1,n_{\mathrm{heads}}=1,d_{\mathrm{model}}=768$) Test accuracy “groks” when training accuracy reaches near-$100\%$. (b) UM: ($m_{\mathrm{test}=2048},\mathrm{depth}=6,n_{\mathrm{heads}}=4,d_{\mathrm{model}}=768$) Test accuracy “groks” simultaneously with training accuracy on sequences with period longer than context length ($\mathcal{T}_{m}>L=256$), indicating delayed discovery of underlying rules.
@@
We then carefully examine the training dynamics in both FM and UM settings (Figure 2). We categorize the training sequences into two groups: i) sequences with periods shorter than the context length, which can be solved through simple copying, and ii) sequences with periods longer than the context length, which require the model to deduce underlying rules for prediction. Our analysis reveals that the model first acquires copying ability for group i) in the early stages of training, and later “groks” the solution for group ii) (Power et al., 2022). Notably, the model’s ability to generalize to test modulus $m_{\mathrm{test}}$ emerges simultaneously with this grokking phenomenon. These results demonstrate that the model develops different capabilities at distinct stages of training, with generalization ability emerging only after the model learns the underlying rules through solving the more challenging sequences. In Appendix E, we present an ablation study where models are trained exclusively on either short-period or long-period sequences. Our findings indicate that training exclusively on long-period sequences enables model generalization.
@@ 4 Interpreting How Transformers Predict PRNGs
In this section, we uncover the underlying algorithms implemented by the models for both FM and UM cases. While certain details of the algorithms differ in the two cases, they share common properties originating from the underlying LCG structure. We first discuss properties of LCG sequences that will be useful in interpreting model behaviors.
4.1 Residual Number System Representations
@@
where $\{\alpha_{0},\dots,\alpha_{10}\}$ are the binary-valued digits (bits).
A useful property of LCGs with modulus $m=2^{(\cdot)}$ is that each digit in the binary representation has a fixed period along the sequence. As shown in Figure 3, for a sequence of period $\mathcal{T}_{m}=m=2^{11}$, the $w^{th}$ lowest digit has a period of $2^{w}$ (Knuth, 1997). Thus, lower (higher) digits have smaller (larger) periods along LCG sequences. (See Section F.1 for a detailed derivation.) We will see later that trained Transformers have emergent structures that find these binary representations and utilize them to make systematic predictions from the context. Notably, the per-digit period plays and important role in the prediction accuracy of that digit. To understand this, consider the $r$-step iteration of Equation 1:
$x_{t+r}=a^{r}x_{t}+\sum_{i=1}^{r}a^{i-1}c\mod m\;.$ (3)
In this new sequence wherein we skip $r$ steps, the period of each digit $\alpha_{w-1}$ reduces from $2^{w}$ to $2^{w}/\gcd(r,2^{w})$. Consequently, the higher digits become relatively simpler to predict due to reduced periods while some lower digits become trivial to predict due to being constant along this new sequence. We demonstrate this for $m=2048$ in Figure 3 (top panel). The digit-wise periods in the new sequence with $r=2$ are reduced by a factor of 2, while the last digit is simply constant. Higher values of $r=2^{k}$ will lead to even further simplifications of the sequence. Transformers can simplify the task of predicting LCG sequences by utilizing $r$-step iterations from the in-context examples – with longer contexts leading to larger values of $r$. Consequently, the per-digit and overall accuracies improve substantially with context (see Figure 4).
While moduli of the form $m=2^{(\cdot)}$ lead Transformers to find binary representations, similar simplifications in composite moduli require more general representations of the Residual Number System (RNS) (Garner, 1959). RNS represents numbers by their values modulo pairwise coprime factorizations of $m$. Specifically, consider sequences with a composite modulus $m$, which has a prime factorization $m=p_{1}^{w_{1}}p_{2}^{w_{2}}\cdots p_{q}^{w_{q}}$. In this case, we can uniquely represent each number $(x\,\mathrm{mod}\,m)$ as the tuple of residuals $(x\,\mathrm{mod}\,p_{1}^{w_{1}},x\,\mathrm{mod}\,p_{2}^{w_{2}},\dots,x\,\mathrm{mod}\,p_{q}^{w_{q}})$. Analogous to Equation 2, we can further decompose each residual,
$x\;\mathrm{mod}\;p_{j}^{w_{j}}=\alpha_{j,0}\,p_{j}^{0}+\alpha_{j,1}\,p_{j}^{1}+\cdots+\alpha_{j,w_{j}-1}\,p_{j}^{w_{j}-1}$ (4)
where $\alpha_{j,w}\in\{0,1,\dots,p_{j}-1\}$ are base-$p_{j}$ digits. We refer to $\{\alpha_{j,w}\}$ as the “RNS representation” in the remainder of the text. When $\mathcal{T}_{m}=m$, each digit $\alpha_{j,w}$ has a period of $p_{j}^{w}$ (derivation in Section F.2). The $r$ step iteration Equation 3 reduces the period of each digit $\alpha_{j,w}$ from $p_{j}^{w}$ to $p_{j}^{w}/\gcd(r,p_{j}^{w})$. This results in simplification of the prediction task whenever $r=p_{1}^{k_{1}}p_{2}^{k_{2}}\cdots p_{q}^{k_{q}}$. We will see that identifying the RNS representations is a key simplification that the Transformer discovers in learning LCG sequences.